4. Five males with a particular genetic disorder have onc child each. I need the second part answer
nt/PlayerHomework.aspx?homeworkld-512478192&questionld-2&flushed false. MATH 2193 Internet Statisics Spring 2019 Homework: Section 5.2 Homework Score: 0 of 1 pt 5.2.7 Save 40119 (3complete) HW Score: 15.79%, 3 of 19pts Question Heip Five males with a particular genetic disorder have one child each The random variable x is the number of children among the five who inherit the genetic disorder. Determine whether the table describes a probability distribution. If it does, find the mean and standard deviation P(x) 00206 0 1209 0283803332 0.1956 00459 Find the mean of the random variable x Select the correct choice below and, if necessary, fill in the answer box to complete your choice A. μ.21 (Round to one decmal place as needed) O B. The table is not a probability distribution Find the standard deviation of the random variable x Select the correct choice below and, if necessary, fill in the answer box to complete your choice @a. O B. σ-Q(Round to one decimal place as needed ) The table is not a probability distribution (00) More Click to select and enter your answer(s) and then click Check Answer 0 Clear All O Type here to seal
Answer #1
x f(x) xP(x) x2P(x)
0 0.021 0.000 0.000
1 0.121 0.121 0.121
2 0.284 0.568 1.135
3 0.333 1.000 2.999
4 0.196 0.782 3.130
5 0.046 0.230 1.148
total 2.700 8.532
E(x) =μ= ΣxP(x) = 2.7000
E(x2) = Σx2P(x) = 8.5320
Var(x)=σ2 = E(x2)-(E(x))2= 1.2420
std deviation=         σ= √σ2 = 1.1145

from above mean μ =2.700

and std deviation σ=1.1